THE Ã STATES OF THE C_{3}-Ar AND C_{3}-Kr VAN DER WAALS COMPLEXES: FLUORESCENCE POLAR-IZATION AND SATURATION

KAN-SEN CHEN, SHIN-SHIN CHENG, JUN-MEI CHAO, ANTHONY J. MERER, \underline{YEN-CHU HSU}, Institute of Atomic and Molecular Sciences, Academia Sinica, P. O. Box 23-166, Taipei 10617, R. O. C..

The degree of polarization of the laser-induced fluoresence of the Ã−X̃ system of the C_{3}-Ar and C_{3}-Kr van der Waals complexes was determined and has been used to probe the orientation of the transition dipole of C_{3} with respect to the principal axis of each complex. The fluorescence polarization was measured near the 2_{0}^{2−}, 2_{0}^{2+}, 1_{0}^{1} and 2_{0}^{1+}3_{0}^{1} bands of C_{3} under supersonic expansion conditions. The results of polarization measurements of several van der Waals spectral features will be compared with those of the associated R(0) lines of free C_{3}. The effect of partial saturation on the fluorescence polarization measurements will be discussed.